• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Roche

    Log entry time 00:02:56 on August 6,2001

    Entry number 69508

    keyword=evening Shift summary

    up to 17:20 continue taking data on Carbon.

    17:20 -> MCC start the change of energy to 3 passes
    -> cycle quadrupoles and go to 1.60 GeV/c for the two magnets

    18:00 -> R. Michael called : he is going to reboot adaql2 which is unable to mount
    foreing disk and therefore copy data on silo. After that bring rc for the 3
    daq up:
    run #621 #622 tests for the daq
    On right arm the "commentary boxes" are not showing up by them selves
    Bob suggest to bring them interactively :
    on adaql2
    cd to scripts
    type epicsRunStart or epicsRunStop

    19:20-> O. Hansen called to announce that he is going to reboot again adaql2
    (he knows Bob did it one hour ago)
    bring runcontrol up and take cosmics

    23:55 -> beam is almost back
    start beam profile,position... studies